OpenStudy (anonymous):

Given that f is a linear function for which f(-3) = 0 and f(3) = 6 determine f(0)=?

OpenStudy (anonymous):

can you write the equation of a line that passes through the points (-3, 0) and (3,6)?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

this is what the question is asking you to do

OpenStudy (anonymous):

oh ok i didnt understand that

OpenStudy (anonymous):

its only asking for a single number

OpenStudy (anonymous):

when you write the equation of this line, plug in 0 for x and you will get your y-coordinate

OpenStudy (anonymous):

yes, but in order to get this number, you must write a linear function that relates the first two values

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

so doing y1-y2/x1-x2 I will get the cord. and the answer

OpenStudy (anonymous):

you will get the slope of the line

OpenStudy (anonymous):

so making it short 6/6?

OpenStudy (anonymous):

yes, now you have the slope

OpenStudy (anonymous):

to get the equation of the line, you need to find the y-intercept

OpenStudy (anonymous):

the y-intercept is the answer to your question in this case

OpenStudy (anonymous):

so then I do y-y1=m(x-x1)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ok and "M" being the 6/6 right?

OpenStudy (anonymous):

yes, you can use 1 for m because 6/6=1

OpenStudy (anonymous):

so would x be the 0 or x1?

OpenStudy (anonymous):

never mind I got it